Comment pquest SNF 2009-08-21 15:39:16: Update heater

eenriquez at snf.stanford.edu eenriquez at snf.stanford.edu
Fri Aug 21 15:39:16 PDT 2009


Completed chamber reassembly.  Ran O2 plasma for 10 minutes with the heater set at 45 C.  The temp reading was stable even when the plasma was on.  Loaded and unloaded wafers with no problem.  The He cooling flow however was around 5.6 sccm.
Need to adjust the wafer clamp to improve the He leak-by.  Also need to re-install all the RF enclosure cover screws




More information about the pquest-pcs mailing list